OpenStudy (anonymous):

Ryan made two similar glass display boxes shaped like triangular prisms. The length of the sides of the base of the first display box is 2 feet, and the length of the sides of the base of the second display box is 6 feet. If the volume of the first display box is 24 ft3, what is the volume of the second display box? Answer 288 ft3 648 ft3 724 ft3 864 ft3

Directrix (directrix):

648 ft3

OpenStudy (anonymous):

how did you get that? like how did you work it out?

OpenStudy (anonymous):

and thank-you

Directrix (directrix):

The scale factor of the two similar boxes is 2/6 or 1/3. If two solids are similar, the cube of the scale factor = the ratio of the volumes. (1/3)^3 = 24/V where V is the volume of the larger solid. 1/27 = 24/v V = 27*24 = 648

OpenStudy (anonymous):

oh ok thank-you so much!